please answer both parts
1)Quantitative Problem: You are holding a portfolio with the following investments and betas:
Stock Dollar investment Beta
A $300,000 1.35
B 200,000 1.60
C 400,000 0.80
D 100,000 -0.35
Total investment $1,000,000
The market's required return is 11% and the risk-free rate is 4%. What is the portfolio's required return? Do not round intermediate calculations. Round your answer to three decimal places.
____ %


2)An individual has $35,000 invested in a stock with a beta of 0.7 and another $55,000 invested in a stock with a beta of 2.5. If these are the only two investments in her portfolio, what is her portfolio's beta? Do not round intermediate calculations. Round your answer to two decimal places.
_____
Answer #1

HI

As per policy we wil solve only top most question here.

1) To find out portfolio required return, we will find portfolio weighted average beta

Portfolio beta= (300,000/1,000,000)* 1.35 +(200,000/1,000,000)* 1.6 +(400,000/1,000,000)* 0.8 -(100,000/1,000,000)* 0.35

Portfolio beta = 1.01

market required return rm = 11%

risk free rate rf = 4%

As per CAPM

Portfolio return = rf+beta*(rm-rf)

=4 + 1.01*(11-4)

=4 + 1.01*7

= 11.07%

THanks